A gas is collected at 22.0 ˚C and 745.0 mm Hg. When the temperature is changed to 0 ˚C, what is the resulting pressure?


1
Expert's answer
2021-10-07T03:44:06-0400

"T_1 = 22+273 =295 \;K \\

T_2 = 0 + 273 = 273 \;K \\

p_1 = 745.0 \;mmHg \\

\frac{p_1}{T_1} = \frac{p_2}{T_2} \\

p_2 = \frac{p_1T_2}{T_1} \\

= \frac{745.0 \times 273}{295} = 689.44 \;mmHg"